# Information Discovery in e-Commerce

### v1

@@ 6.5.3 Explainable e-commerce recommendation
Although recommendation models can generate relevant items for users in many e-commerce applications, it is often ambiguous to understand why an item is recommended to a user. Hence it is necessary to develop explainable recommendation strategies to generate not only high-quality recommendations but also intuitive explanations. Recent years have witnessed increasing publications on explainable recommendation. zhang2014explicit generated textual sentences as recommendation explanation to help users understand each recommendation result. chen2018visually proposed visually explainable recommendations where particular regions of a recommended image are highlighted as the visual explanations for users. sharma2013social, quijano2017make generated a list of social friends who also liked the recommended product as social explanations for target user, whereas gao2019bloma generated the recommendation described by a set of topics.
Several researchers have started to generate explanations for deep recommendation models. For example, several studies leverage knowledge graph for interpretation. They aim to construct multi-hop paths from users to items along the knowledge graph, which indicates a specific explainable user-item relation [hu2018leveraging, wang2019explainable, xian2019reinforcement]. Besides, chen2021neural proposed a neural collaborative reasoning system integrating the power of representation learning and logical reasoning. However, research on explainable deep recommendation models is relatively new and deserve to be further explored in e-commerce.
@@ 6.5.4 Biases and debiasing in recommendations
Many recommendation solutions about fitting user behaviors may deteriorate owing to biases in behavior inherent in e-commerce recommendation  [LightGCN, sun2019bert4rec]. In e-commerce scenarios, user behaviors are observational rather than experimental, which is often affected by many factors, e.g., self-selection of the user (selection bias) [marlin2007collaborative], systematic exposure mechanisms (exposure bias) [ovaisi2020correcting], public opinions (conformity bias) [krishnan2014methodology, liang2016modeling] and the display position (position bias) [joachims2007evaluating]. These biases make the data deviate from reflecting true preferences of users in recommender systems. Efforts to debias recommendation can be divided into three major categories: (i) data imputation, which assigns pseudo-labels to missing data to reduce variance [steck2013evaluation], (ii) inverse propensity scoring (IPS), which reweighs the collected data for an expectation-unbiased learning [sun2019debiasing, wang2016learning], and (iii) generative modeling, which assumes the generation process of data and reduces biases [liang2016modeling]. Most approaches lack the universal capacity to account for mixed or even unknown biases. To bridge the gap, chen2021autodebias proposed a universal debiasing framework that not only account for multiple biases and their combinations, but also frees human efforts to identify biases. 10.1145/3488560.3498375 introduced DANCER, a debiasing method that accounts for dynamic selection bias and user preferences, demonstrating its improved rating prediction performance over static bias methods. 10.1145/3583780.3615011 explored the use of uncertainty estimates in ranking scores to reduce societal biases in retrieved documents while minimizing utility loss. They proposed an uncertainty-aware, post hoc bias mitigation method that outperforms existing baselines in terms of utility-fairness trade-offs, controllability, and computational costs, without requiring additional training. Although recent years have seen a surge in research effort on recommendation biases, biases are still a huge problem in e-commerce recommender systems. Sophisticated meta models to capture complex patterns and exploration of dynamic biases in recommendation should provide helpful insights.
@@ 6.5.5 Unifying recommendation and search
Search and recommendation in e-commerce have similar characteristics, except for the different representation of “contexts” — search aims at retrieving relevant items for matching “query” while recommendation aims at finding items for matching user preference. However, researchers usually conduct separate studies on them and use different techniques and training data for the two tasks. Thus, building a unified model for search and recommendation has the potential to improve both tasks as more comprehensive user behavior data can be used. One practical way to unify the two tasks is aforementioned conversational recommendation, and the other is personalized search, which we detail next.
@@
Early search engines, like Google and AltaVista, retrieved personalized results based on keywords. Personalized search, as pioneered by Google, has become far more complex with the goal to “understand exactly what you mean and give you exactly what you want.” Concretely, a personalized search engine not only focuses on retrieving items that satisfy the user’s current information needs, which is usually related to the query topic, but also considers user personality and aims at retrieving items that meet user preference. To achieve both goals, it is critical to model the interactions between users, items and queries. ai2017learning leveraged hierarchical embedding model to linearly combine the item-query matching scores with item-user preference scores; guo2019attentive explored long and short term user preference learning model for personalized search; 10.1145/3459637.3482489 integrated user behaviors in search and recommendation into a heterogeneous behavior sequence and used a joint model to handle both tasks based on this unified sequence; si2023search leveraged users’ search interests for recommendations. It separately learns similar and dissimilar representations from search and recommendation behaviors using transformer encoders; liu2020structural constructed a specific user-item-query graph and conducts node representation learning on the graph. 10.1145/3488560.3498414 proposed a method that jointly predict user clicks for both search and recommendation scenarios by constructing a unified graph to share user and item representations uniformly. Such kind of graph embedding techniques open the potential to integrate both node information and topological structure information, which can capture high-order user-item-query interactions.
6.5.6 LLMs in recommendation
@@
Large language models (LLMs) have exhibited strong capabilities in understanding and processing text and researchers hope to apply LLMs to recommendation systems. The main benefit of using LLMs in recommendation systems is their ability to produce high-quality representations of text features and make use of the wide range of knowledge they hold [liu2023pre]. LLM-based models can capture context more accurately, allowing them to understand better user questions, product descriptions, and other textual information. Existing studies to apply LLMs to recommendation systems can be divided into two categories: discriminative strategies and generative strategies.
@@
For studies about discriminative strategies, to improve the quality of vector representations for queries and products, and fully utilize the external knowledge stored in LLMs, a common approach is to fine-tune the original models, adapting them to recommendation tasks in order to obtain high-quality representations. qiu2021u proposed a novel U-BERT approach that utilizes a pre-training and fine-tuning framework to learn user representations. By leveraging content-rich domains, U-BERT compensates for users’ features in domains where behavior data is insufficient, improving recommendation performance. Similarly,  wu2021userbert utilized unlabeled user behavior data and incorporates two self-supervised tasks: masked behavior prediction and behavior sequence matching for user model training.
@@
Compared to discriminative models, generative models have better natural language generation capabilities. Therefore, most generative models typically translate recommendation tasks into natural language tasks, allowing the model to directly output recommendation results through fine-tuning or in-context learning. sun2023chatgpt introduced a sliding window prompt strategy for ranking candidates. This strategy ranks items within a window at each step, sliding the window from back to front multiple times to generate the final ranking results. This approach helps improve ranking performance by iteratively refining the candidate list. kang2023llms investigated the ability of LLMs to predict user ratings based on past behavior, comparing them with traditional Collaborative Filtering (CF) methods.
Given existing advances for such area, there is still challenges which guide the future research for LLMs in e-commerce recommender systems
@@
Chapter 3 provides insights on how natural language processing technologies have been widely applied in e-commerce platform interfaces to help consumers better communicate with e-commerce portals. This chapter demonstrates question answering (QA) services and dialogue systems in e-commerce portals. We divide this chapter into three sections: e-commerce question answering, e-commerce dialogue systems, and emerging directions. Specifically, we first detail characteristics and existing approaches to e-commerce question answering (Section 7.1). Then, we demonstrate recent studies on dialogue systems applicable in e-commerce customer services (Section 7.2). Lastly, we briefly describe emerging directions in e-commerce question answering and dialogue systems (Section 7.3).
@@ 7.1 Question answering in e-commerce
In this section, we describe related work on e-commerce question answering. We divide this section into $3$ parts: we first briefly introduce existing studies on question answering in Section 7.1.1, then in Section 7.1.2 we formulate characteristics of product-aware question answering; eventually, we detail existing approaches on e-commerce question answering in Section 7.1.3 and Section 7.1.4.
@@ 7.1.1 Introduction to question answering
We first provide a brief introduction to question answering (QA) research. QA systems [simmons1965answering] greatly facilitate users’ access to information for many web-based applications, where QA services provide a proper answer to a given question from the user [heilman2010tree, li2002learning]. Question answering research has received much attention in the past decades, including approaches to question classification, answer selection, answer generation, and answer summarization [li2002learning, heilman2010tree, liu2016retrieving, geigle2016scaling, song2017summarizing]. QA systems have various classifications. According to the dimensionality of questions, existing QA systems can be divided into open-domain and specific-domain QA systems [chen2020open]. Open-domain QA focuses on answering questions relying on knowledge and ontologies [ferrucci2010building], whereas specific-domain QA focuses on providing proper answers in a specific scenario, e.g., customer service, hotel booking, etc. Existing QA systems can be divided into retrieval-based and generation-based QA systems according to how they generate answers [yang2015wikiqa]. The former searches and extracts potential answers via search engines, whereas the latter applies generation-based methods to give proper answers to the questions. According to the answers, QA systems can be divided into factoid QA and non-factoid QA [song2017summarizing]. Factoid QA systems return a concise answer to the given question, whereas non-factoid QA systems provide more subjective answers to the given questions.
@@
During the early stage of QA studies, there are $4$ categories of QA systems: list-structured database systems, graphical database systems, text-based systems, and logical inference systems forms [simmons1965answering]. All these systems focus on the limited scope with rule-based strategies. Search engines remain integral components of QA systems. With the development of information retrieval, researchers find that QA systems can be a new research direction [jurafsky2000speech]. In 1999, TREC-QA tracks were proposed to provide benchmarks for researchers [srihari1999information, voorhees1999trec]. There are $3$ main components in a retrieval-based TREC-QA system: question processing, passage retrieval, and answer processing [jurafsky2000speech]. For each step, sub-tasks must be considered, e.g., query formulation or answer type detection. Based on such a framework, several approaches have been proposed to address research tasks in each component [brill2001data, li2002learning]. Early studies on QA focus on factoid QA systems that generate concise answers [srihari1999information, jurafsky2000speech, brill2001data]. Retrieval-based methods effectively answer these concise and simple questions [jurafsky2000speech, ahn2004using]. However, complicated questions are found difficult to be addressed by pure retrieval-based methods [lin2006role]. Therefore, integrating natural language understanding and knowledge reasoning techniques is essential for retrieval-based QA strategies in answering complicated questions.
@@
In TREC-QA 2004, questions are grouped into topics, which motivates research on fact identification from reference knowledge resources, e.g., Wikipedia [ahn2004using]. Wikipedia can be considered a generic collection of articles with real-world facts for open-domain QA systems. With the development of knowledge bases (KBs), innovations have occurred in the context of QA from KBs with the creation of resources like web questions and short questions [berant2013semantic, bordes2015large]. However, inherent limitations such as incompleteness and fixed schemas still exist in traditional KB-based QA systems. Thus increasing QA studies generate answers from raw text explored on Wikipedia [ryu2014open, ahn2004using, buscaldi2006mining, ferrucci2010building]. As far as we know, ahn2004using first combined Wikipedia as a text resource with another. Similarly, ryu2014open performed a QA system using a Wikipedia-based knowledge model by combining articles with other answer-matching components. ferrucci2010building, baudivs2015yodaqa integrated web-based and Wikipedia-based articles as knowledge resources into highly developed full-pipeline QA platforms.
@@
In recent years, increasing QA models have been proposed by applying deep neural networks to understand the questions and generate answers. [yin2015neural] presented an end-to-end neural network model, namely neural generative question answering (GENQA), that can generate answers to simple factoid questions. A bi-directional attention flow mechanism was then proposed to obtain a query-aware passage representation [seo2016bidirectional]. chen2017reading developed a system for question answering from Wikipedia, namely DrQA, composed of a two-stage retrieval-reader QA framework. DrQA includes a document retriever module via bigram hashing and TF-IDF matching. It also contains a document reader module where a multi-layer recurrent neural network is trained to detect answer spans in those few returned documents.
@@
Following chen2017reading, most existing open-domain QA systems apply a two-stage retrieval-reader framework in their QA mechanisms [wang2018r, sun2018open, lin2018denoising, pang2019has, lee2019latent, guu2020retrieval, karpukhin2020dense, izacard2021leveraging, mao2020generation, sachan2021end, NEURIPS2021_da3fde15, yu-etal-2022-kg, DBLP:conf/emnlp/KediaZL22, DBLP:conf/emnlp/Ju00Z022, DBLP:conf/acl/WangY023]. These studies employed a determinate retrieval function and treated each passage independently in the retrieval stage: wang2018r proposed a reinforcement learning-based ranking strategy in the retrieval stage. sun2018open offered a graph convolution-based neural network by operating over heterogeneous graphs of KB facts and text sentences. In contrast with previous KB-based open-domain QA systems, the authors propose heterogeneous update rules that handle KB nodes differently from the text nodes. lin2018denoising proposed a coarse-to-fine denoising model to extract correct answers from multiple paragraphs in the noisy data. Their model employed a paragraph selector to filter out those noisy paragraphs and keep those informative paragraphs. Similarly, pang2019has proposed a three-level probabilistic formulation model for open-domain QA. Word-level matching strategies are usually applied in the retrieval stage to match keywords represented in high-dimensional and sparse vectors. Dense passage retrieval has been successfully applied to open-domain QA to improve the matching performance as it is complementary to sparse representations in the retrieval stage. karpukhin2020dense aimed to train a dense embedding model using only pairs of questions and passages. izacard2021leveraging proposed an effective two-step dense passage retrieval method. First, the authors retrieved supporting passages using either sparse or dense embeddings, then employed a sequence-to-sequence model to generate the answer. zhu2021adaptive used the partially observed Markov decision process (POMDP) to formulate the QA research problem to address the mismatch problem through the determinate retrieval function. The authors applied a reinforcement learning method to reflect the interactive characteristics between the QA model (i.e., agent), including a belief module with a set of evidence for state performance, and the intractable large-scale corpus. yu-etal-2022-kg leveraged a knowledge graph to establish relational dependencies among retrieved passages and employed a graph neural network to re-rank retrieved passages for each query. Recently, several studies have been proposed to improve reader performance and thereby improve QA performance. DBLP:conf/emnlp/KediaZL22 proposed a method fusing information across multiple passages within a transformer encoder using global representation tokens. DBLP:conf/emnlp/Ju00Z022 proposed a knowledge graph enhanced passage reader which fuses graph and contextual representations into the hidden states of the reader model. DBLP:conf/acl/WangY023 enhanced the Fusion-in-Decoder (FiD) framework by incorporating a rationale identification process to distinguish between relevant and spurious passages, thereby improving the model’s reasoning and performance in open-domain QA.
@@
Relevant studies have also been proposed to enhance QA performance through reading comprehension. A model that matches the question with a passage using gated attention-based recurrent networks has been verified effective in QA benchmark datasets [wang2017gated]. QANet was proposed to combine local convolution with global self-attention for reading comprehension, which improved the reading comprehension performances [yu2018qanet]. Recent studies also verified that pre-trained language models effectively understand questions and answers in QA systems [guu2020retrieval, mao2020generation, sachan2021end, NEURIPS2021_da3fde15]. mao2020generation augmented a query in open-domain QA using text generation of a pre-trained language model. sachan2021end proposed a QA method with an unsupervised pre-training of the retriever with a supervised fine-tune procedure.
@@
During past decades, many benchmark QA datasets have been proposed. Several QA benchmark datasets, such as SQuAD [rajpurkar2016squad], TriviaQA [joshi2017triviaqa], and SearchQA [dunn2017searchqa], only evaluate the reasoning ability within a single paragraph, whereas the other relevant documents or paragraphs are neglected. Moreover, these datasets employ existing knowledge bases for multi-hop reasoning, constrained by the schema of existing knowledge bases. yang2018hotpotqa proposed an open-domain QA benchmark dataset, namely HotpotQA, which requires reasoning over multiple documents without constraining itself to an existing knowledge base. To understand how the questions and answers are distributed in open-domain QA, lewis2020question performed large-scale analysis on open-domain QA benchmark datasets, and provided annotated subsets of test sets indicating whether test-time questions are duplicates of training time questions.

### v3

@@ 6.5.3 Explainable e-commerce recommendation
Although recommendation models can generate relevant items for users in many e-commerce applications, it is often ambiguous to understand why an item is recommended to a user. Hence it is necessary to develop explainable recommendation strategies to generate not only high-quality recommendations but also intuitive explanations. Recent years have witnessed a growth in the number of publications on explainable recommendation. zhang2014explicit generated textual sentences as recommendation explanation to help users understand each recommendation result. chen2018visually propose visually explainable recommendations where particular regions of a recommended image are highlighted as the visual explanations for users. sharma2013social, quijano2017make generate a list of social friends who also like the recommended product as social explanations for target user, whereas gao2019bloma generate the recommendation described by a set of topics.
Several researchers have started to generate explanations for deep recommendation models. For example, several studies use knowledge graphs for interpretation. They construct multi-hop paths from users to items along the knowledge graph, which indicates a specific explainable user-item relation [hu2018leveraging, wang2019explainable, xian2019reinforcement]. Besides, chen2021neural propose a neural collaborative reasoning system integrating the power of representation learning and logical reasoning. However, research on explainable deep recommendation models is relatively new and deserves to be further explored in e-commerce.
@@ 6.5.4 Biases and debiasing in recommendations
Many recommendation solutions about fitting user behavior may deteriorate owing to biases in behavior inherent in e-commerce recommendation  [LightGCN, sun2019bert4rec]. In e-commerce scenarios, user behavior is observational rather than experimental, which is often affected by many factors, e.g., self-selection of the user (selection bias) [marlin2007collaborative], systematic exposure mechanisms (exposure bias) [ovaisi2020correcting], public opinions (conformity bias) [krishnan2014methodology, liang2016modeling] and the display position (position bias) [joachims2007evaluating]. These biases make the data deviate from reflecting true preferences of users in recommender systems. Efforts to debias recommendation can be divided into three major categories: (i) data imputation, which assigns pseudo-labels to missing data to reduce variance [steck2013evaluation], (ii) inverse propensity scoring (IPS), which reweighs the collected data for an expectation-unbiased learning [sun2019debiasing, wang2016learning], and (iii) generative modeling, which assumes the generation process of data and reduces biases [liang2016modeling]. Most approaches lack the universal capacity to account for mixed or even unknown biases. To bridge the gap, chen2021autodebias propose a universal debiasing framework that not only account for multiple biases and their combinations, but also frees human efforts to identify biases. huang2022clickrec introduce DANCER, a debiasing method that accounts for dynamic selection bias and user preferences, demonstrating its improved rating prediction performance over static bias methods. 10.1145/3583780.3615011 explore the use of uncertainty estimates in ranking scores to reduce societal biases in retrieved documents while minimizing utility loss. They propose an uncertainty-aware, post hoc bias mitigation method that outperforms baselines in terms of utility-fairness trade-offs, controllability, and computational costs, without requiring additional training. Although recent years have seen a surge in research efforts devoted to recommendation biases, biases are still an important problem in e-commerce recommender systems. Sophisticated meta models to capture complex patterns and exploration of dynamic biases in recommendation should provide helpful insights.
@@ 6.5.5 Unifying recommendation and search
Search and recommendation in e-commerce have similar characteristics, except for the different representation of “contexts” – search aims at retrieving relevant items for matching a query while recommendation aims at finding items for matching a user’s preferences. However, researchers usually conduct separate studies on them and use different techniques and training data for the two tasks. Thus, building a unified model for search and recommendation has the potential to improve both tasks as more comprehensive user behavior data can be used. One practical way to unify the two tasks is as part of the aforementioned conversational recommendation scenario, and the other is personalized search, which we detail next.
@@
Early search engines, like Google and AltaVista, retrieved personalized results based on keywords. Personalized search has become far more complex with the goal to “understand exactly what you mean and give you exactly what you want.” Concretely, a personalized search engine not only focuses on retrieving items that satisfy the user’s current information needs, which is usually related to the query topic, but also considers user personality and aims at retrieving items that meet user preference. To achieve both goals, it is critical to model interactions between users, items and queries. ai2017learning use a hierarchical embedding model to linearly combine the item-query matching scores with item-user preference scores; guo2019attentive explore long and short term user preference learning model for personalized search; 10.1145/3459637.3482489 integrate user behavior in search and recommendation into a heterogeneous behavior sequence and use a joint model to handle both tasks based on this unified sequence; si2023search use users’ search interests for recommendations; they separately learns similar and dissimilar representations from search and recommendation behaviors using transformer encoders. liu2020structural construct a specific user-item-query graph and conducts node representation learning on the graph. 10.1145/3488560.3498414 propose a method that jointly predict user clicks for both search and recommendation scenarios by constructing a unified graph to share user and item representations uniformly. Such graph embedding techniques open the potential to integrate both node information and topological structure information, which can capture high-order user-item-query interactions.
6.5.6 Large language models in recommendation
@@
Large language models (LLMs) have exhibited strong capabilities in understanding and processing text. Their application to recommendation systems is actively being explored. The main benefit of using LLMs in recommendation systems is their ability to produce high-quality representations of text features and make use of the wide range of knowledge they hold [liu2023pre]. LLM-based models can capture context more accurately, allowing them to better understand user questions, product descriptions, and other textual information. Studies that apply LLMs to recommendation systems can be divided into two categories: discriminative strategies and generative strategies.
@@
For studies into discriminative strategies, to improve the quality of vector representations for queries and products, and fully use the external knowledge stored in LLMs, a common approach is to fine-tune the original models, adapting them to recommendation tasks in order to obtain high-quality representations. qiu2021u propose a novel U-BERT approach that utilizes a pre-training and fine-tuning framework to learn user representations. By using content-rich domains, U-BERT compensates for users’ features in domains where behavior data is insufficient, improving recommendation performance. Similarly, wu2021userbert use unlabeled user behavior data and incorporate two self-supervised tasks: masked behavior prediction and behavior sequence matching for user model training.
@@
Compared to discriminative models, generative models have better natural language generation capabilities. Therefore, most generative models typically translate recommendation tasks into natural language tasks, allowing the model to directly output recommendation results through fine-tuning or in-context learning. sun2023chatgpt introduce a sliding window prompt strategy for ranking candidates. This strategy ranks items within a window at each step, sliding the window from back to front multiple times to generate the final ranking results. This approach helps improve ranking performance by iteratively refining the candidate list. kang2023llms investigate the ability of LLMs to predict user ratings based on past behavior, comparing them with traditional collaborative filtering methods.
The application of LLMs to e-commerce recommender systems is still in its early days. Many challenges remain, in terms of evaluation, effectiveness, efficiency, and transparency.
@@
Section 3 provides insights on how natural language processing technologies have been widely applied in e-commerce platform interfaces to help consumers better communicate with those platforms. This section zooms in on question answering (QA) services and dialogue systems on e-commerce platforms. We divide this section into three parts: e-commerce question answering, e-commerce dialogue systems, and emerging directions. We first detail characteristics and approaches to e-commerce question answering (Section 7.1). Then, we demonstrate recent studies on dialogue systems applicable in e-commerce customer services (Section 7.2). Lastly, we describe emerging directions in e-commerce question answering and dialogue systems (Section 7.3).
@@ 7.1 Question answering in e-commerce
In this section, we describe related work on e-commerce question answering. We divide this section into three parts: we first introduce studies on question answering in Section 7.1.1, then in Section 7.1.2 we formulate characteristics of product-aware question answering; finally, we detail approaches to e-commerce question answering in Section 7.1.3 and 7.1.4.
@@ 7.1.1 Introduction to question answering
QA systems [simmons1965answering] are meant to facilitate users’ access to information. For many web-based applications QA services provide a proper answer to a given question from the user [heilman2010tree, li2002learning]. Question answering research has received much attention in the past decades, including approaches to question classification, answer selection, answer generation, and answer summarization [li2002learning, heilman2010tree, liu2016retrieving, geigle2016scaling, song2017summarizing]. QA systems have various classifications. QA systems can be divided into open-domain and domain-specific QA systems [chen2020open]. Open-domain QA focuses on answering questions relying on knowledge and ontologies [ferrucci2010building], whereas domain-specific QA focuses on providing proper answers in a specific scenario, e.g., customer service, hotel booking, etc. QA systems can also be divided into retrieval-based and generation-based QA systems according to how they generate answers [yang2015wikiqa]. The former searches and extracts potential answers via search engines, whereas the latter applies generation-based methods to give proper answers to the questions. And finally, according to the answers, QA systems can be divided into factoid QA and non-factoid QA [song2017summarizing]. Factoid QA systems return a concise answer to the given question, whereas non-factoid QA systems provide more subjective answers to the given questions.
@@
Early work on QA distinguishes between four categories of QA system: list-structured database systems, graphical database systems, text-based systems, and logical inference systems [simmons1965answering]. All these systems have a limited scope with their rule-based strategies. Search engines remain integral components of QA systems. With the development of information retrieval, research “re-discovered” QA systems in the late 1990s [jurafsky2000speech]. TREC has launched dedicated QA tracks in 1999, with the purpose of advancing research into QA systems [srihari1999information, voorhees1999trec]. A typical retrieval-based TREC QA system has three main components: question processing, passage retrieval, and answer processing [jurafsky2000speech]. For each step, sub-tasks must be considered, e.g., query formulation or answer type detection. Based on this framework, several approaches have been proposed to address research tasks in each component [brill2001data, li2002learning]. Early studies on QA focus on factoid QA systems that generate concise answers [srihari1999information, jurafsky2000speech, brill2001data]. Retrieval-based methods effectively answer these concise and simple questions [jurafsky2000speech, ahn2004using]. However, complicated questions are found difficult to be addressed by pure retrieval-based methods [lin2006role]. Therefore, integrating natural language understanding and knowledge-based reasoning techniques is essential for retrieval-based QA strategies in answering complicated questions.
@@
In TREC-QA 2004, questions are grouped into topics, which motivates research on fact identification from reference knowledge resources, e.g., Wikipedia [ahn2004using]. Wikipedia can be considered a generic collection of articles with real-world facts for open-domain QA systems. With the development of knowledge bases, innovations have occurred in the context of QA from knowledge bases with the creation of resources like web questions and short questions [berant2013semantic, bordes2015large]. However, inherent limitations such as incompleteness and fixed schemas have persisted in traditional knowledge-based QA systems. Thus, in the 2000s QA work increasingly on systems that are able to generate answers from raw text explored, especially using Wikipedia [ahn2004using, buscaldi2006mining, ferrucci2010building, ryu2014open]. As far as we know, ahn2004using are the first to combine Wikipedia as a text resource with other resources in QA. Similarly, ryu2014open perform QA using a Wikipedia-based knowledge model by combining articles with other answer-matching components. ferrucci2010building, baudivs2015yodaqa integrate web-based and Wikipedia-based articles as knowledge resources into highly developed full-pipeline QA platforms.
@@
In more recent years, QA models increasingly apply deep neural networks to understand questions and generate answers. yin2015neural present an end-to-end neural network model, neural generative question answering (GENQA), that can generate answers to simple factoid questions. Subsequently, a bi-directional attention flow mechanism has been proposed to obtain query-aware passage representations [seo2016bidirectional]. chen2017reading develop a system for question answering from Wikipedia, DrQA, that is composed of a two-stage retrieval-reader QA framework. DrQA includes a document retriever module based on bigram hashing and TF-IDF matching. It also contains a document reader module where a multi-layer recurrent neural network is trained to detect answer spans in those few returned documents.
@@
Following chen2017reading, most open-domain QA systems apply a two-stage retrieval-reader framework in their QA mechanisms [wang2018r, sun2018open, lin2018denoising, pang2019has, lee2019latent, guu2020retrieval, karpukhin2020dense, izacard2021leveraging, mao2020generation, sachan2021end, NEURIPS2021_da3fde15, yu-etal-2022-kg, DBLP:conf/emnlp/KediaZL22, DBLP:conf/emnlp/Ju00Z022, DBLP:conf/acl/WangY023]. These studies employ a determinate retrieval function and treat each passage independently in the retrieval stage. wang2018r propose a reinforcement learning-based ranking strategy in the retrieval stage. sun2018open offer a graph convolution-based neural network by operating over heterogeneous graphs of knowledge base facts and text sentences. In contrast with previous kowledge-based open-domain QA systems, the authors propose heterogeneous update rules that handle knowledge base nodes differently from the text nodes. lin2018denoising design a coarse-to-fine denoising model to extract correct answers from multiple paragraphs in the noisy data. Their model employs a paragraph selector to filter out those noisy paragraphs and keep informative paragraphs. Similarly, pang2019has describe a three-level probabilistic formulation model for open-domain QA. Word-level matching strategies are usually applied in the retrieval stage to match keywords represented in high-dimensional and sparse vectors. Dense passage retrieval has successfully been applied to open-domain QA to improve the matching performance as it is complementary to sparse representations in the retrieval stage. karpukhin2020dense train a dense embedding model using only pairs of questions and passages. izacard2021leveraging detail an effective two-step dense passage retrieval method; the authors retrieve supporting passages using either sparse or dense embeddings and then employ a sequence-to-sequence model to generate the answer. zhu2021adaptive use a partially observed Markov decision process (POMDP) to re-formulate the QA problem using a reinforcement learning method to optimize the interactions between different components. yu-etal-2022-kg use a knowledge graph to establish relational dependencies among retrieved passages and employ a graph neural network to re-rank retrieved passages for each query. More recent work has proposed to improve reader performance and thereby improve QA performance. DBLP:conf/emnlp/KediaZL22 introduce a method for fusing information across multiple passages within a transformer encoder using global representation tokens. DBLP:conf/emnlp/Ju00Z022 design a knowledge graph enhanced passage reader that fuses graph and contextual representations into the hidden states of the reader model. DBLP:conf/acl/WangY023 enhance the fusion-in-decoder (FiD) framework by incorporating a process to distinguish between relevant and spurious passages, thereby improving the model’s reasoning and performance in open-domain QA.
@@
A model that matches the question with a passage using gated attention-based recurrent networks has been shown to be effective on QA benchmark datasets [wang2017gated]. QANet combines local convolution with global self-attention for reading comprehension, which improved the reading comprehension performances [yu2018qanet]. More recent studies have shown that pre-trained language models effectively understand questions and answers in QA systems [guu2020retrieval, mao2020generation, sachan2021end, NEURIPS2021_da3fde15]. mao2020generation augment a query in open-domain QA using text generation of a pre-trained language model. sachan2021end propose a QA method with an unsupervised pre-training of the retriever with a supervised fine-tune procedure.
@@
Many benchmark QA datasets have been proposed. Several QA benchmark datasets, such as SQuAD [rajpurkar2016squad], TriviaQA [joshi2017triviaqa], and SearchQA [dunn2017searchqa], only evaluate the reasoning ability within a single paragraph, whereas the other relevant documents or paragraphs are neglected. These datasets employ knowledge bases for multi-hop reasoning, and are therefore constrained by the schema of knowledge bases. yang2018hotpotqa introduce an open-domain QA benchmark dataset, HotpotQA, which requires reasoning over multiple documents without constraining itself to a knowledge base. To understand how the questions and answers are distributed in open-domain QA, lewis2020question perform a large-scale analysis on open-domain QA benchmark datasets, and provide annotated subsets of test sets indicating whether test-time questions are duplicates of training time questions.